OpenStudy (anonymous):

solve this system of equations by elimination 5x^2-5y^2=-35 5x^2+2y^2=77 please show how to solve problem

OpenStudy (unklerhaukus):

\[5x^2-5y^2=-35\\ 5x^2+2y^2=77\] what do you get when you find the difference between the lefthand sides, and righthand sides? \[(5x^2-5y^2)-[5x^2+2y^2]=(-35)-[77]\] Hopefully the\(x\)- terms will disappear, and you'll have an equation that you can solve for \(y\)